OpenStudy (anonymous):

which ordered pair is the solution of this equation -3x - 7y=37 a. -3,-4 b. -4,-4 c. -4,-3 d. 4,4

OpenStudy (anonymous):

a

OpenStudy (anonymous):

-3(-3)-7(-4) = 9 + 28 =37

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

the answer is a.

OpenStudy (anonymous):

-3,-4

OpenStudy (anonymous):

okhay thanks do you want a medal

OpenStudy (anonymous):

if u like

OpenStudy (anonymous):

ok i gave you one

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i'll be your fan

OpenStudy (anonymous):

ok